OpenStudy (moderator):

Find the values of x in the following equation: -2|-5x + 2| + 2 = -74

OpenStudy (anonymous):

add 2 divide by -2 then remove absolute values signs

OpenStudy (moderator):

satellite u have team viewer its a free download and u can help people ??

OpenStudy (anonymous):

\[-2|-5x+2|+2=-74\] \[-2|-5x+2|=-72\] \[|-5x+2|=36\]

OpenStudy (anonymous):

then solve \[5x-2=36\] and \[5x-2=-36\] go get your answers.

OpenStudy (anonymous):

@moderator i have no idea what team viewer is

OpenStudy (moderator):

its like a free program u download and u can like get on their computer and help them

OpenStudy (anonymous):

\[5x-2=36\] \[5x=38\] \[x=\frac{38}{5}\] \[5x-2=-36\] \[5x=-34\] \[x=-\frac{34}{5}\]

OpenStudy (anonymous):

help whom?

OpenStudy (moderator):

u cud help me a little

OpenStudy (moderator):

and itss free